Solve Hermite's probabilistic equation given by L[u]=u''-xu=-nu
Answer #1

SOLUTION:

From given data,

Solve Hermite's probabilistic equation given by L[u]=u''-xu'=-nu

Let , L[u]=u''-x u'= - nu = > L[u]=u''-x u'+nu = 0

L[u]=u''-x u'+nu = 0 -------- (1)

Let ,

ke

kak.r

u'' = \sum_{k=0}^{\infty } k(k-1) a_{k}x^{k-2}

By substitution above all in equation (1)

\sum_{k=0}^{\infty } k(k-1) a_{k}x^{k-2}-kak1+n\sum_{k=0}^{\infty } a_{k}x^{k}=0

\sum_{k=0}^{\infty } k(k-1) a_{k}x^{k-2}- x\sum_{k=0}^{\infty } k a_{k}x^{k}Σ nakrk

By replace

k = k-2

\sum_{k=0}^{\infty } (k+2)*(k+1) * a_{k+2} * x^{k}+\sum_{k=0}^{\infty } (n-k)* a_{k}*x^{k}=0

\sum_{k=0}^{\infty }[ (k+1) *(k+2)* a_{k+2}+(n-k)a_{k}]x^{k} = 0

Where,

a_{k+2} = (k-n)*ak / (k+1)(k+2)

Let us consider the recursion formula:

At k=0

a_{2} = - n*a0 / 1*2

At k=1

a_{3} = (1-n)*a1 / (1+1)(1+2)

a_{3} =(1-n)*a1 / 2*3

At k=2

a_{4} = [ (2-n) /3*4 ] * [(-n) /2*1] * a0

At k=3

a_{5} = [ (3-n) /4*5 ] * [(1-n) /2*3] * a1

Also ,

ke

= (a0 + a1x + a2 x2 + ...........)

= (a0 - (na0 / 2*1)* x2...........) + (a1 x - ((1-n)a1 / 2*3)* x3 +...........)

= (a0 - (na0 / 2*1)* x2 - (n(2-n) / 4*3*2*1)* x4 +...........) + (a1 x - ((1-n)a1 / 2*3)* x3 + ((3-n)(1-n) / 5*4*3*2)* x5...........)

= (a0*(1-n / 2*1)* x2 - (n(2-n) / 4*3*2*1)* x4+..........) + (a1 *(x - ((1-n)/ 2*3)* x3 + ((n-1)(n-3) / 5*4*3*2)* x5...........)

= (a0*(1- n / 2!)* x2 - (n(2-n) / 4!)* x4+..........) + (a1 *(x - ((n-1)/ 3!)* x3 + ((n-1)(n-3) / 5!)* x5...........)

This is the solution

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.